thestrugglesreal:

What is the median for the data set? 574, 526, 512, 579, 595, 517, 524, 552, 558, 541 Express your answer as a decimal to the nearest tenth.

ThisGirlPretty:

Hello! Welcome to QuestionCove! :)! Do you know how to find the median? @thestrugglesreal

thestrugglesreal:

HI ! and yes the middle number

thestrugglesreal:

595, 517

ThisGirlPretty:

Give me one second please.

ShadowWolf:

I think 546.5

ShadowWolf:

yup 546.5 is the right answer

ThisGirlPretty:

Can you explain how you got it so the user would understand how to get it please? :)

thestrugglesreal:

What is the interquartile range for the data set? 241, 230, 201, 245, 209, 211, 242, 201, 204, 228, 242, 243

thestrugglesreal:

35.5 or 44?

thestrugglesreal:

@ThisGirlPretty

thestrugglesreal:

@ShadowWolf

ShadowWolf:

I think the interquartile range is 44.

thestrugglesreal:

you sure?

ShadowWolf:

I am not sure

ShadowWolf:

@Hero

thestrugglesreal:

hmm

ShadowWolf:

@Vocaloid

thestrugglesreal:

hmm im not so sure about that

thestrugglesreal:

anyone ?

thestrugglesreal:

not that either

ShadowWolf:

I think it would be 44

thestrugglesreal:

ima go with 35.5

ShadowWolf:

ok Hope u pass

Hero:

Explain how either of you two got those answers. Do you really know how to solve that? If so, how about explaining your steps instead of just posting answers.

Hero:

I want to see if anyone really knows what they're doing here.

thestrugglesreal:

i just got both of them from searching the question up. (brainly/PA) but some said 35.5 and others said 44, sooo that's why i posted it here

Hero:

smh. So that means you don't really know how to do it.

thestrugglesreal:

correct. x'D

Hero:

Here's how to do it. 1st put the data set in numerical order: Then split it up in half: 1st half is Q1 | 2nd half is Q2: Q1 | 201, 201, 204, 209, 211, 228 Q2 | 230, 241, 242, 245, 245, 243 Next find the median of Q1 and Q2 Median of Q1 = (204 + 209)/2 = 206.5 Median of Q2 = (242 + 245)/2 = 243.5 Finally subtract: M(Q2) - M(Q1) = 243.5 - 206.5 = 37 Therefore, the interquartile range for the given data set is 37.

Hero:

After doing some research, I have discovered that there are various methods for calculating interquartile range depending on your level.

daDreamersHERE:

I don't know if it helps now. I do not know the answer to the problem but I do know how to find interquartile range. I learned that you first find the median of that data set. Then, you will have numbers on both sides of the median. The left side of the median is quartile 1 and the right side of the median is quartile 3. You find the median of both quartiles. Last, you see the difference between the two quartiles once you find the difference between those two quartiles. That's the interquartile range.